# Shortcuts to Adiabaticity via Adaptive Quantum Zeno Measurements

### v1

@@ Abstract
CD techniques rely on coherent quantum control, although they can be generalized to open quantum systems with nonunitary dynamics [41, 2]. Other prominent approaches to control quantum evolution utilize quantum measurements. An instance is the celebrated quantum Zeno effect, which relies on a sequence of projective measurements, and has long been recognized as a mechanism to freeze the time evolution and the coupling to the surrounding environment [31, 28, 18]. The evolution of a system with Hamiltonian $H$, subject to frequent projective measurements associated with a projector $P$, is described by the effective Hamiltonian $H_{\rm Z}=PHP$ in the Zeno limit [20, 17]. Recently, the quantum Zeno effect was used experimentally to guide the evolution of a quantum system [23]. Such quantum Zeno dragging is achieved by continuous quantum measurements of a time-varying observable. The analogy between this approach and CD has been pointed out [30].
@@
$P(t_{k+1})e^{-iH\delta t}P(t_{k})=P-i\left(PHP-i[\dot{P},P]\right)\delta t.$ (3)
@@
where the non-Hermitian term $\Gamma_{\kappa}(t)$ quantifies the leakage induced by finite absorption strength. As before, if $P(t)$ is chosen to be a spectral projector of $H(t)$ (or a sum of spectral projectors), the projected term satisfies $P(t)H(t)P(t)=H(t)$ within the corresponding eigenspace, and the geometric term $i[\dot{P}(t),P(t)]$ coincides with the gauge-invariant CD Hamiltonian that enforces adiabatic transport. Thus, a strong time-dependent absorbing potential provides a dynamical realization of transitionless evolution that is formally equivalent, at leading order, to both stroboscopic Zeno measurements and CD. In the static case, Schulman established a relation between the Zeno dynamics induced by stroboscopic measurements and complex absorbing potentials [35, 19, 32], which was experimentally verified [37]. Such a connection was established by analyzing the response time of a two-level system, leading to a relation between $\delta t$ and $\kappa$. As shown in the End Matter, such a relation can be generalized by comparing the generators of evolution, rather than the response time scale. Indeed, this can be done in a model-independent way for time-dependent Zeno measurements. Specifically, it can be shown that
@@
Appendix A Finite-time-step corrections to Zeno dragging
@@
up to $\mathcal{O}(\delta t^{3})$ corrections. Equation (A) makes explicit that leakage is generically $\mathcal{O}(\delta t^{2})$ for a state initially confined to $\mathrm{Ran}P(t)$.
To connect with the Zeno Hamiltonian, it is convenient to factor the one-step map into a unitary part generated by the ideal Zeno Hamiltonian plus a small nonunitary correction. Using the effective Zeno Hamiltonian $H_{\rm Z}(t)\equiv P(t)H(t)P(t)+i[\dot{P}(t),P(t)]$, one finds that the expansion (A) can be reorganized as
@@
whose expectation value controls the norm loss (A). Equivalently, the conditioned evolution within the Zeno subspace can be described by an effective non-Hermitian time-dependent Hamiltonian,
@@
For a time-independent projector ($\dot{P}=0$), Eq. (A) reduces to the standard static-Zeno leakage $p_{\rm leak}\simeq\delta t^{2}\langle HQH\rangle$ [17]. For spectral projectors $P(t)$ of $H(t)$, the projected term $PHP$ equals the corresponding energy block, and the geometric term in $H_{\rm Z}$ coincides with the CD connection. Eq. (A) then isolates the finite $\delta t$ error due to incomplete suppression of off-block transitions.
Appendix B Relation between stroboscopic measurements and complex absorbing potentials

### v2

@@ Abstract
CD techniques rely on coherent quantum control, although they can be generalized to open quantum systems with nonunitary dynamics [43, 2]. Other prominent approaches to control quantum evolution utilize quantum measurements. An instance is the celebrated quantum Zeno effect, which relies on a sequence of projective measurements, and has long been recognized as a mechanism to freeze the time evolution and the coupling to the surrounding environment [33, 29, 19]. The evolution of a system with Hamiltonian $H$, subject to frequent projective measurements associated with a projector $P$, is described by the effective Hamiltonian $H_{\rm Z}=PHP$ in the Zeno limit [21, 18]. Recently, the quantum Zeno effect was used experimentally to guide the evolution of a quantum system [24]. Such quantum Zeno dragging is achieved by continuous quantum measurements of a time-varying observable. The analogy between this approach and CD has been pointed out [32]. Adaptive Zeno measurements have also been proposed to engineer quantum heat engines [8] and for ground state preparation [31, 46].
@@
$P(t_{k+1})e^{-iH\delta t}P(t_{k})=\left[\mathbb{I}-i\delta t\left(PHP-i[\dot{P},P]\right)\right]P.$ (3)
@@
where the non-Hermitian term $\Gamma_{\kappa}(t)$ quantifies the leakage induced by finite absorption strength. As before, if $P(t)$ is chosen to be a spectral projector of $H(t)$ (or a sum of spectral projectors), the projected term satisfies $P(t)H(t)P(t)=H(t)$ within the corresponding eigenspace, and the geometric term $i[\dot{P}(t),P(t)]$ coincides with the gauge-invariant CD Hamiltonian that enforces adiabatic transport, as detailed in the End Matter. Thus, a strong time-dependent absorbing potential provides a dynamical realization of transitionless evolution that is formally equivalent, at leading order, to both stroboscopic Zeno measurements and CD. In the static case, Schulman established a relation between the Zeno dynamics induced by stroboscopic measurements and complex absorbing potentials [37, 20, 34], which was experimentally verified [39]. Such a connection was established by analyzing the response time of a two-level system, leading to a relation between $\delta t$ and $\kappa$. As shown in the End Matter, such a relation can be generalized by comparing the generators of evolution, rather than the response time scale. Indeed, this can be done in a model-independent way for time-dependent Zeno measurements. Specifically, it can be shown that
@@
Appendix A Operational implementation of nonselective stroboscopic PVM measurements
In the stroboscopic construction, we repeatedly apply the completely positive map
$\rho\mapsto\sum_{n}P_{n}(t_{k})\rho P_{n}(t_{k}),$ (26)
corresponding to a full projection-valued measurement (PVM) $\{P_{n}(t_{k})\}$. This map does not require averaging over different experimental runs. It can be implemented on a single quantum system by coupling the system to an ancillary probe (Stinespring dilation), measuring the probe, and resetting it between steps.
Let the probe Hilbert space admit an orthonormal pointer basis $\{|n\rangle_{A}\}$. Prepare the probe in a reference state $|0\rangle_{A}$ and apply a unitary interaction of the form
$U_{k}=\sum_{n}P_{n}(t_{k})\otimes W_{n},\qquad W_{n}|0\rangle_{A}=|n\rangle_{A},$ (27)
which entangles the system sectors with orthogonal probe states. Acting on an arbitrary system state $\rho$, this unitary produces
$U_{k}\big(\rho\otimes|0\rangle\langle 0|\big)U_{k}^{\dagger}=\sum_{n,m}P_{n}(t_{k})\,\rho\,P_{m}(t_{k})\otimes|n\rangle\langle m|.$ (28)
A projective measurement of the probe in the pointer basis $\{|n\rangle_{A}\}$ yields a definite outcome $n_{k}$ and collapses the system to $P_{n_{k}}(t_{k})\rho P_{n_{k}}(t_{k})/p_{n_{k}}$. The probe is then reset to $|0\rangle_{A}$, and the evolution proceeds to the next stroboscopic interval. However, if the measurement outcome is not used to condition subsequent control operations (nonselective protocol), the reduced state of the system after the probe measurement is obtained by tracing over the probe and results in Eq. (26). This implies discarding the classical measurement record while retaining the same physical system. No restart of the experiment is required and the system continues evolving within whichever sector was realized in that particular run.
Appendix B Finite-time-step corrections to Zeno dragging
@@
up to $\mathcal{O}(\delta t^{3})$ corrections. Equation (B) makes explicit that leakage is generically $\mathcal{O}(\delta t^{2})$ for a state initially confined to $\mathrm{Ran}P(t)$.
To connect with the Zeno Hamiltonian, it is convenient to factor the one-step map into a unitary part generated by the ideal Zeno Hamiltonian plus a small nonunitary correction. Using the effective Zeno Hamiltonian $H_{\rm Z}(t)\equiv P(t)H(t)P(t)+i[\dot{P}(t),P(t)]$, one finds that the expansion (B) can be reorganized as
@@
whose expectation value controls the norm loss (B). Equivalently, the conditioned evolution within the Zeno subspace can be described by an effective non-Hermitian time-dependent Hamiltonian,
@@
For a time-independent projector ($\dot{P}=0$), Eq. (B) reduces to the standard static-Zeno leakage $p_{\rm leak}\simeq\delta t^{2}\langle HQH\rangle$ [18]. For spectral projectors $P(t)$ of $H(t)$, the projected term $PHP$ equals the corresponding energy block, and the geometric term in $H_{\rm Z}$ coincides with the CD connection. Eq. (B) then isolates the finite $\delta t$ error due to incomplete suppression of off-block transitions.
Appendix C Multi-Sector Complex Absorbing Potentials
Consider the non-Hermitian Hamiltonian
$H_{\rm nh}(t)=H(t)-i\kappa\,\Lambda(t),\qquad\Lambda(t)=\sum_{n}\lambda_{n}P_{n}(t),$ (36)
with real coefficients $\lambda_{n}$ and $\kappa>0$ large. If all $\lambda_{n}$ are equal, then $\Lambda(t)=\lambda\mathbb{I}$ and it induces only a global decay; therefore nontrivial sector separation requires $\lambda_{n}\neq\lambda_{m}$ for some $n\neq m$.
In the co-moving frame $|\tilde{\psi}\rangle=W^{\dagger}|\psi\rangle$, the Schrödinger equation becomes
$i\partial_{t}|\tilde{\psi}\rangle=\big(\tilde{H}(t)-i\kappa\Lambda_{0}\big)|\tilde{\psi}\rangle,\qquad\Lambda_{0}=\sum_{n}\lambda_{n}P_{n}(0),$ (37)
with $\tilde{H}(t)=W^{\dagger}HW+iW^{\dagger}\dot{W}$. Decomposing $|\tilde{\psi}\rangle=\sum_{n}P_{n}(0)|\tilde{\psi}\rangle\equiv\sum_{n}|\tilde{\psi}_{n}\rangle$ yields
$i\partial_{t}|\tilde{\psi}_{n}\rangle=\sum_{m}\tilde{H}_{nm}(t)|\tilde{\psi}_{m}\rangle-i\kappa\lambda_{n}|\tilde{\psi}_{n}\rangle,$ (38)
where $\tilde{H}_{nm}=P_{n}(0)\tilde{H}P_{m}(0)$. For $n\neq m$, transitions are suppressed when $\kappa|\lambda_{n}-\lambda_{m}|\gg\|\tilde{H}_{nm}(t)\|.$ Under this condition, adiabatic elimination of off-diagonal sectors gives a leading-order effective generator
$H_{\rm eff}(t)=H_{Z}(t)-i\kappa\sum_{n}\lambda_{n}P_{n}(t)+\mathcal{O}(\kappa^{-1}),$ (39)
which is block diagonal in the monitored decomposition.
A special case is the two-sector decomposition defined by a single projector $P(t)$ and its complement $Q(t)=\mathbb{I}-P(t)$. Choosing $\lambda_{P}=0$, $\lambda_{Q}=1$, one recovers $H_{\rm nh}(t)=H(t)-i\kappa Q(t)$, discussed in the main text. In this case, the effective Zeno Hamiltonian becomes $H_{\rm eff}(t)=P(t)H(t)P(t)+i[\dot{P}(t),P(t)]+\mathcal{O}(\kappa^{-1})$, demonstrating that the multi-sector construction reduces to the standard Kato–Avron Hamiltonian for a protected subspace.
Appendix D Relation between stroboscopic measurements and complex absorbing potentials